Find the radius of a curve banked at an angle of 35 degrees negotiated by a car at a velocity of 20.0 m/s?


1
Expert's answer
2021-11-29T11:44:42-0500

The radius is

"r=\frac{v^2}{g\tan\theta}=\frac{20^2}{9.8\tan35°}=58.3\text{ m}."